OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5/9 (F − 32), where F stands for degrees Fahrenheit. Solve the formula for F. Show all work. Please help!!

OpenStudy (anonymous):

C = 5/9(F - 32) C = 5/9F - 160/9 (multiply by 9 to get rid of fractions) 9C = 5F - 160 9C + 160 = 5F (9C + 160)/5 = F You might need to get a second opinion on this.....I am not 100% sure about this. You wouldn't happen to have answer choices, would you ?

OpenStudy (anonymous):

@bahrom7893 .....can you check this please

OpenStudy (anonymous):

would you go a step farther and say 9/5C + 32 = F ??

OpenStudy (anonymous):

or am I totally off on this question ?

OpenStudy (anonymous):

@kelliegirl33 no sorry there isnt any answer choices..but i'll try you're answer and hope for the best i guess!

OpenStudy (anonymous):

@kelliegirl33 Thanks for your answer though!

OpenStudy (anonymous):

wait...I don't know if it would be (9C + 160)/5 = F ,or 9/5C + 32 = F ???

OpenStudy (anonymous):

I would go with (9C + 160)/5

OpenStudy (anonymous):

@kelliegirl33 Alright. Thanks!

OpenStudy (anonymous):

I hope it is correct....just to let you know, if this was for my grade, I would put that answer